In a simple random sample of 95 families, 70 had one or more pets at home. At the 1% level of significance, test the claim that the proportion of families with one or more pets differs from 60%.

Calculate the test statistic, round your answer to two decimal places.
Answer #1

Given :

Sample size = n = 95

Number of success = X = 70

Sample proportion = {\hat{p}} = X/n = 70/95 = 0.7368

Hypothesis test :

The null and alternative hypothesis is

Ho : p = 0.60

Ha : p \neq 0.60

Test statistic :

z = ({\hat{p}} - p)/[√p(1-p)/n]

= (0.7368 - 0.60)/[√0.60(1-0.60)/95]

= 2.72

Therefore the test statistic is z = 2.72...(if getting wrong try 2.79)